OpenStudy (anonymous):

If you had x+5 over -x Would your answer be -x+5/x?

OpenStudy (sidsiddhartha):

\[\frac{ x }{ -x }=-1\]

OpenStudy (sidsiddhartha):

\[\frac{ 5 }{ -x }=-\frac{ 5}{ x }\]

OpenStudy (mathstudent55):

\(\dfrac{x + 5}{-x} = \dfrac{x + 5}{-1 \cdot x} = -1 \cdot \dfrac{x + 5}{ x} = \dfrac{-1 \cdot (x + 5)}{ x} =\dfrac{-x - 5}{x} \)

OpenStudy (mathstudent55):

The -1 you factor out of the denominator has to be multiplied by the numerator. The numerator becomes -(x + 5) which after distributing the -1, then becomes -x - 5.